The temperature of an ideal gas was held at 400 degrees C. The initial pressure and volume were 750 torr and 50 ml respectively. What would the final pressure be if the volume were increased to 1500 ml?
Answer by venugopalramana(3286)

The temperature of an ideal gas was held at 400 degrees C. The initial pressure and volume were 750 torr and 50 ml respectively. What would the final pressure be if the volume were increased to 1500 ml?
AT CONSTANT TEMPERATURE,FOR AN IDEAL GAS PRODUCT OF TOTAL PRESSURE AND VOLUME IS CONSTANT....
P1=750 TORR.....V1=50 ML.
P2=?............V2=1500 ML
P1V1=P2V2
P2=P1V1/V2=750*50/1500=25 TORR